One number is 5 greater than another. The product of the number is 36. Find the number
Answer by unlockmath(1688)

Hello,
Let's show the number as n. The equation will look like this:
n(n+5)=36
Multiply it out.
n^2+5n=36
Subtract 36 to get:
n^2+5n-36=0
Factor:
(n+9)(n-4)=0
Solve for n
n=-9
n=4
Actually both numbers work. Check it out.
Make sense?
RJ
www.math-unlock.com